A fielder can throw a cricket ball to maximum horizontal distance of 100m. How high the fielder can throw the same ball?
A. 25m
B. 50m
C. 40m
D. 100m

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
We know that horizontal range is maximum if the angle of projection is 45o and is given by
Rmax = u2 / g
Here, Rmax = 100 m
so, u2 / g  = 100 m
If the cricketer throws the ball vertically upwardthen the ball will attain the maximum height from the ground
Hmax = u2 / 2g  = 100 /2 = 50 m.